# Orientable and negative orientable sequences

### v1

@@ Abstract
Analogously to de Bruijn sequences, orientable sequences have application in automatic position-location applications and, until recently, studies of these sequences focused on the binary case. In recent work by Alhakim et al., a range of methods of construction were described for orientable sequences over arbitrary finite alphabets; some of these methods involve using negative orientable sequences as a building block. In this paper we describe three techniques for generating such negative orientable sequences, as well as upper bounds on their period. We then go on to show how these negative orientable sequences can be used to generate orientable sequences with period close to the maximum possible for every non-binary alphabet size and for every tuple length. In doing so we use two closely related approaches described by Alhakim et al.
@@ 1 Introduction
Orientable sequences, i.e. periodic sequences with elements drawn from a finite alphabet with the property that any subsequence of $n$ consecutive elements (an $n$-tuple) occurs at most once in either direction, were introduced in 1992 [3, 4, 6]. They are of interest due to their application in certain position-resolution scenarios. For the binary case, a construction and an upper bound on the period were established by Dai et al. [6], and further constructions were established by Mitchell and Wild [12] and Gabrić and Sawada [9]. A bound on the period and methods of construction for $q$-ary alphabet sequences (for arbitrary $q$) were given by Alhakim et al. [2].
In this paper we use two of the methods proposed in [2] to construct $q$-ary orientable sequences with near-optimal period for arbitrary $n$ and $q$. We first demonstrate methods for generating negative orientable sequences for near-optimal periods for every $q>2$ and $n$. These can in turn be used to generate orientable sequences via the inverse Lempel Homomorphism, following [2].
This parallels recent work by Gabrić and Sawada [9] who also constructed large-period $q$-ary orientable sequences for arbitrary $n$ and $q$, albeit using a different approach.
@@ 1.1 Basic terminology
An $n$-window sequence $S=(s_{i})$ is said to be an orientable sequence of order $n$ (an $\mathcal{OS}_{q}(n)$) if $s_{n}(i)\neq s_{n}(j)^{R}$, for any $i,j$.
We also need two related definitions.
@@
An $n$-window sequence $S=(s_{i})$ is said to be a negative orientable sequence of order $n$ (a $\mathcal{NOS}_{q}(n)$) if $\mathbf{s}_{n}(i)\not=-{\mathbf{s}_{n}(j)}^{R}$, for any $i,j$.
As discussed in Alhalkim at al. [2], it turns out that negative orientable sequences are of importance in constructing orientable sequences. Observe that a sequence is orientable if and only if it is negative orientable for the case $q=2$. Also note that if $S=(s_{i})$ is orientable or negative orientable then so is its negative ($-s_{i}$).
1.2 The de Bruijn graph and the Lempel Homomorphism
Definition 1.4
For a nonzero $\beta\in\mathbb{Z}_{q}$, we define a function $D_{\beta}$ from $B_{n}(q)$ to $B_{n-1}(q)$ as follows. For $a=(a_{1},\dots,a_{n})$ and $b=(b_{1},\dots,b_{n-1}),\ D_{\beta}(a)=b$ if and only if $b_{i}=d_{\beta}(a_{i},a_{i+1})$ for $i=1$ to $n-1$, where $d_{\beta}(a_{i},a_{i+1})=\beta(a_{i+1}-a_{i})\mod q.$
$D((s_{i}))=\{(t_{i}):t_{j}=\beta(s_{j+1}-s_{j})\}.$
We are particularly interested in the case $\beta=1$, and we simply write $D$ for $D_{1}$. The weight $w(S)$ of a cycle $S$ is the weight of the ring sequence corresponding to $S$ (that is the sum of the terms $s_{0},s_{1},\dots,s_{m-1}$ treating $s_{i}$ as an integer in the range $[0,q-1]$). Similarly we write $w_{q}(S)$ for $w(S)\bmod{q}$.
@@ 2 Constructing maximal orientable sequences
We start by showing that, unlike the binary case, orientable sequences for $n=2$ with period meeting the bound of [2] can easily be constructed for every $q>2$. Observe that analogous results have recently been obtained by Gabrić and Sawada [8].
@@
There are simple algorithms for finding Eulerian circuits — see for example Gibbons [10, Figure 6.5]. Moreover, a direct construction of an $OS_{q}(2)$ for the case where $q>2$ is a prime is presented in [2, Construction 5.3].
@@ 3 Negative orientable sequences
We next establish some fundamental results on negative orientable sequences (see Definition 1.3), given their importance in constructing orientable sequences using the methods of [2]. We further describe three methods of construction for such sequences.
@@ 3.1 A bound on the period
We start by giving a bound on the period of negative orientable sequences.
A $q$-ary $n$-tuple $\mathbf{u}=(u_{0},u_{1},\dots,u_{n-1})$ is said to be negasymmetric if $u_{i}=-u_{n-1-i}$ for every $i$ ($10\leq i\leq n-$ 1). Clearly an $\mathcal{NOS}_{q}(n)$ cannot contain any negasymmetric $n$-tuples. This motivates the following.
@@
    The $n$ even case is immediate. For $n$ odd ($n=2m+1$ say), an $n$-tuple $(u_{0},u_{1},\ldots,u_{m-1},u_{m},-u_{m-1},\ldots,-u_{1},-u_{0})$ is negasymmetric if and only if $u_{m}=-u_{m}$. If $q$ is odd this implies $u_{m}=0$, and if $q$ is even then this implies $u_{m}$ is either 00 or $m$. The result follows.
@@
    There are trivially $q^{n}$ $q$-ary $n$-tuples, of which only those that are non-negasymmetric can occur in a $\mathcal{NOS}_{q}(n)$. If $\mathbf{u}$ is such an $n$-tuple, then at most one of $\mathbf{u}$ and $-\mathbf{u}$ can occur in an $\mathcal{NOS}_{q}(n)$. The result follows.
@@
Clearly an $n$-tuple is $n$-negasymmetric if and only is it is negasymmetric. We also need the notion of uniformity.
@@
That is, $L(\mathbf{v})$ is simply the set of $n$-tuples whose first $n-1$ entries equal $\mathbf{v}$. Clearly, the sets $L(\mathbf{v})$ for all $(n-1)$-tuples $\mathbf{v}$ are disjoint. We are interested in how the sets $L(\mathbf{v})$ intersect with the set of $n$-tuples occurring in either $S$ or $S^{R}$ when $S$ is a $\mathcal{NOS}_{q}(n)$ and $\mathbf{v}$ is a negasymmetric $(n-1)$-tuple.
@@
$L^{*}_{S}(\mathbf{v})=\{\mathbf{u}\in L(\mathbf{v}):~{}\mathbf{u}~{}\mbox{appears in}~{}S~{}\mbox{or}~{}-S^{R}\}.$
It is important to observe that, from the definition of negative orientable, an element of $L^{*}_{S}(\mathbf{v})$ (as with any $n$-tuple) can only appear at most once in $S$ or $-S^{R}$. We now state a result on which our bound is built.
@@
    Suppose $L^{*}_{S}(\mathbf{v})$ is non-empty. Then, by definition, and (without loss of generality) assuming that an element of $L(\mathbf{v})$ occurs in $S$ (as opposed to $S^{R}$), we know that $v_{i}=s_{j+i}$ for some $j$ ($20\leq i\leq n-$ 2). Since $\mathbf{v}$ is negasymmetric, it follows immediately that $v_{i}=-s_{j+n-2-i}$ ($20\leq i\leq n-$ 2). That is, for each occurrence of an element of $L(\mathbf{v})$ in $S$, there is an occurrence of a (necessarily distinct) element of $L(\mathbf{v})$ in $-S^{R}$, and vice versa. The result follows.
That is, if $|L(\mathbf{v})|$ is odd, this shows that $S$ and $-S^{R}$ combined must omit at least one of the $n$-tuples in $L(\mathbf{v})$. We can now state our main result.
@@
  • First suppose that $q$ is odd. Suppose also that $\mathbf{v}$ is a negasymmetric non-uniform $(n-1)$-tuple. From Lemma 3.1 there are $q^{\lfloor(n-1)/2\rfloor}$ negasymmetric $(n-1)$-tuples of which precisely one, namely the all-zero tuple, is uniform (since $q$ is odd); i.e. there are $q^{\lfloor(n-1)/2\rfloor}-1$ such $\mathbf{v}$. From Lemma 3.5, none of the $q$ elements of $L(\mathbf{v})$ are negasymmetric, and hence $|L(\mathbf{v})\cap N^{*}_{q}(n)|=q$. Since $|L(\mathbf{v})|=q$ is odd, Lemma 3.9 implies that $L^{*}_{S}(\mathbf{v})$ must omit at least one of the $n$-tuples in $L(\mathbf{v})$. That is, there are $q^{\lfloor(n-1)/2\rfloor}-1$ non-negasymmetric $n$-tuples which cannot appear in $S$ or $-S$. The bound for $q$ odd follows from Corollary 3.2.
@@ 3.2 Construction I: Sequences with $n=2$
Consider the directed graph $G_{q}$ with vertex set $\mathbb{Z}_{q}$, $q\geq 3$, and all directed arcs $(x,y)$, $x,y\in\mathbb{Z}_{q}$, except for $(x,-x)$, $x\in\mathbb{Z}_{q}$. Then a $\mathcal{NOS}_{q}(2)$ $S$ corresponds to a circuit in $G_{q}$. Moreover $-S^{R}$ is also such a circuit and these two circuits are arc-disjoint. We say that a circuit $C=[c_{0},c_{1},\dots,c_{m-1}]$ of length $m$, consisting of arcs $(c_{i},c_{i+1})$, $i=0,\dots,m-1$, identifying $c_{m}$ with $c_{0}$, in $G_{q}$ is good if it is arc-disjoint from $-C^{R}$. We note that if $C^{\prime}$ is another good circuit in $G_{q}$ that is arc-disjoint from $C$ and $-C^{R}$ and if it shares a vertex with $C$ then $C$ and $C^{\prime}$ may be joined to make a good circuit of length the sum of the lengths of $C$ and $C^{\prime}$.
@@
$C_{0}=[0110220\dots 0\frac{q-1}{2}\frac{q-1}{2}]$ of length $3\frac{q-1}{2}$
$C_{1}=[121(-2)131(-3)1\dots 1\frac{q-1}{2}1(-\frac{q-1}{2})]$ of length $4\frac{q-3}{2}$
@@
$C_{i}=[i(i+1)i-(i+1)i\dots i\frac{q-1}{2}i-\frac{q-1}{2}]$ of length $4\frac{q-i-2}{2}$
@@
$C_{\frac{q-3}{2}}=[\frac{q-3}{2}\frac{q-1}{2}\frac{q-3}{2}-\frac{q-1}{2}]$ of length $4$.
These circuits are arc-disjoint since by construction the arcs of $C_{i}$ involve the vertex $i$, with the exception of the loops $(j,j)$ in $C_{0}$, in such a way that there is no repetition. Moreover they are arc-disjoint to $-C_{i}^{R}$, $i=0,\dots,\frac{q-3}{2}$ since by construction the arcs of $-C_{i}^{R}$ involve the vertices $-j$ in such a way that there is no repetition. Thus they are mutually good circuits and it may be seen that these circuits may be joined to give a $\mathcal{NOS}_{q}(2)$ $C$. A simple calculation shows that $C$ has period $\frac{q(q-1)}{2}$ and therefore is maximal by Theorem 3.10.
@@
$C_{0}=[0110220\dots 0\frac{q-2}{2}\frac{q-2}{2}]$ of length $3\frac{q-2}{2}$
$C_{1}=[121(-2)131(-3)1\dots 1\frac{q-2}{2}1(-\frac{q-2}{2})1\frac{q}{2}]$ of length $24\frac{q-4}{2}+$ 2
@@
$C_{i}=[i(i+1)i-(i+1)i\dots i\frac{q-2}{2}i-\frac{q-2}{2}i\frac{q}{2}]$ of length $24\frac{q-2i-2}{2}+$ 2
@@
$C_{\frac{q-4}{)}{2}}=[\frac{q-4}{2}\frac{q-2}{2}\frac{q-4}{2}-\frac{q-2}{2}\frac{q-4}{2}\frac{q}{2}]$ of length $6$
$C_{\frac{q-2}{2}}=[\frac{q-2}{2}\frac{q}{2}]$ of length $2$.

### v5

@@ Abstract
Analogously to de Bruijn sequences, orientable sequences have application in automatic position-location applications and, until recently, studies of these sequences focused on the binary case. In recent work by Alhakim et al., a range of methods of construction were described for orientable sequences over arbitrary finite alphabets; some of these methods involve using negative orientable sequences as a building block. In this paper we describe three techniques for generating such negative orientable sequences, as well as upper bounds on their period. We then go on to show how these negative orientable sequences can be used to generate orientable sequences for every non-binary alphabet size and for every tuple length. In doing so we use two closely related approaches described by Alhakim et al. The periods of both negative orientable and orientable sequences that we construct are of the same order of magnitude as the upper bounds.
@@ 1 Introduction
Orientable sequences are periodic sequences with elements drawn from a finite alphabet with the property that any subsequence of $n$ consecutive elements (an $n$-tuple) occurs at most once in either direction; they were introduced in 1992 [3, 4, 6]. For example, the sequence over $\{0,1,2\}$ with period $[012]$ is orientable as the $2$-tuples it contains are $(0,1),(1,2),(2,0)$ in the forward direction and $(0,2),(2,1),(1,0)$ in the reverse direction. They are of interest due to their application in certain position-resolution scenarios. For the binary case, a construction and an upper bound on the period were established by Dai et al. [6], and further constructions were established by Mitchell and Wild [13] and Gabrić and Sawada [8, 9]. A bound on the period and methods of construction for $q$-ary alphabet sequences (for arbitrary $q$) were given by Alhakim et al. [2].
In this paper we use two of the methods proposed in [2] to construct $q$-ary orientable sequences for arbitrary $n$ and $q$. This provides orientable sequences with period of the same order of magnitude as the upper bound. Both these methods involve the use of $q$-ary negative orientable sequences as a building block. Negative orientable sequences, first introduced in [2], are analogous to orientable sequences except that any $n$-tuple occurs at most once in either a period of the sequence or a period of the negative of the reverse sequence. Interpreting the alphabet of the example above as residues modulo $3$, the example above is not negative orientable as the negatives of the reverse $2$-tuples are $(0,1),(1,2),(2,0)$, equal to the forward $2$-tuples. The sequence, over the same alphabet, with period $[011]$ is negative orientable (but not orientable) as the forward $2$-tuples are $(0,1),(1,1),(1,0)$ and the negatives of the reverse $2$-tuples are $(2,0),(2,2),(0,2)$.
We first demonstrate methods for generating negative orientable sequences for every $q>2$ and $n$. All these sequences can in turn be used to generate orientable sequences via the inverse Lempel Homomorphism, following [2]. This parallels recent work by Gabrić and Sawada [8] who also constructed $q$-ary orientable sequences for arbitrary $n$ and $q$, albeit using a different approach.
The main motivation in studying negative orientable sequences has been to help construct new examples of orientable sequences. However, they may have applications of their own, and potentially merit further study.
The remainder of this paper is structured as follows. In the rest of Section 1, basic concepts are introduced. Section 2 describes how maximal length orientable sequences can be constructed for every alphabet size $q>2$ when the window size is 2. Section 3 establishes an upper bound on the period of negative orientable sequences and gives three methods of construction for such sequences. In Sections 4 and 5, we show how the negative orientable sequences constructed in Section 3 can be used to construct new orientable sequences. Finally, Section 6 concludes the paper, and suggests directions for future work.
@@ 1.1 Basic terminology
A $q$-ary $n$-window sequence $S=(s_{i})$ is said to be an orientable sequence of order $n$ (an $\mathcal{OS}_{q}(n)$) if $s_{n}(i)\neq s_{n}(j)^{R}$, for any $i,j$.
A range of examples of orientable sequences can be found in Gabrić and Sawada [8]. We also need two related definitions.
@@
A $q$-ary $n$-window sequence $S=(s_{i})$ is said to be a negative orientable sequence of order $n$ (a $\mathcal{NOS}_{q}(n)$) if $\mathbf{s}_{n}(i)\not=-{\mathbf{s}_{n}(j)}^{R}$, for any $i,j$.
Examples of negative orientable sequences are given in Section 3.5. As discussed in Alhalkim et al. [2], it turns out that negative orientable sequences can be used to construct orientable sequences. Observe that a sequence is orientable if and only if it is negative orientable for the case $q=2$. Also note that if $S=(s_{i})$ is orientable or negative orientable then so is its negative ($-s_{i}$).
@@ 2 Constructing maximal orientable sequences
We start by showing that, unlike the binary case, orientable sequences for $n=2$ with period meeting the bound of [2] can easily be constructed for every $q>2$. Analogous results have recently been obtained by Gabrić and Sawada [10].
The bound of Theorem 4.11 of [2] is as follows: Suppose that $S=(s_{i})$ is an $\mathcal{OS}_{q}(n)$ ($q\geq 2$, $n\geq 2$). Then the period of $S$ is at most
$(q^{n}-q^{\lceil n/2\rceil}-q^{\lceil(n-1)/2\rceil}+q)/2$ if q is $\mbox{if $q$ is odd},$
$(q^{n}-q^{\lceil n/2\rceil}-q)/2$ if q is $\mbox{if $q$ is even}.$
Further, if $q$ is odd and $n\geq 6$ then the period of $S$ is at most
$(q^{n}-2q^{n/2}-q^{(n-2)/2}+2q)/2$ if n is $\mbox{if $n$ is even},$
$(q^{n}-q^{(n+1)/2}-2q^{(n-1)/2}+q+q^{2})/2$ if n is $\mbox{if $n$ is odd}.$
We note that all these bounds are asymptotically of the same order as $q^{n}/2$.
@@
There are simple algorithms for finding Eulerian circuits — see for example Gibbons [11, Figure 6.5]. Moreover, a direct construction of an $OS_{q}(2)$ for the case where $q>2$ is a prime is presented in [2, Construction 5.3].
@@ 3 Negative orientable sequences
We next establish some fundamental results on negative orientable sequences (see Definition 1.3), given their importance in constructing orientable sequences using the methods of [2]. We further describe three methods of construction for such sequences. Here and throughout the remainder of the paper we only consider the case $q>2$ for two main reasons. As observed in Section 1, when $q=2$, negative orientable sequences are the same as orientable sequences. Further, the methods of construction we present in this section do not work in the case $q=2$.
@@ 3.1 A bound on the period
We start by giving a bound on the period of negative orientable sequences. To establish this bound we need to introduce the following terminology to distinguish a special class of $n$-tuples. A $q$-ary $n$-tuple $\mathbf{u}=(u_{0},u_{1},\dots,u_{n-1})$ is said to be negasymmetric if $u_{i}=-u_{n-1-i}$ for every $i$ ($10\leq i\leq n-$ 1), i.e. if $\mathbf{u}=-\mathbf{u}^{R}$. Clearly an $\mathcal{NOS}_{q}(n)$ cannot contain any negasymmetric $n$-tuples. This motivates the following.
@@
    The $n$ even case is immediate. For $n$ odd ($n=2m+1$ say), an $n$-tuple $(u_{0},u_{1},\ldots,u_{m-1},u_{m},-u_{m-1},\ldots,-u_{1},-u_{0})$ is negasymmetric if and only if $u_{m}=-u_{m}$. If $q$ is odd this implies $u_{m}=0$, and if $q$ is even then this implies $u_{m}$ is either 00 or $q/2$. The result follows.
@@
    There are trivially $q^{n}$ $q$-ary $n$-tuples, of which only those that are non-negasymmetric can occur in a $\mathcal{NOS}_{q}(n)$. If $\mathbf{u}$ is such an $n$-tuple, then at most one of $\mathbf{u}$ and $-\mathbf{u}^{R}$ can occur in an $\mathcal{NOS}_{q}(n)$. The result follows.
@@
Clearly an $n$-tuple is $n$-negasymmetric if and only if is it is negasymmetric. We also need the notion of uniformity.
@@
That is, $L(\mathbf{v})$ is simply the set of $n$-tuples whose first $n-1$ entries equal $\mathbf{v}$. Clearly, the sets $L(\mathbf{v})$ for all $(n-1)$-tuples $\mathbf{v}$ are disjoint. We are interested in how the sets $L(\mathbf{v})$ intersect with the set of $n$-tuples occurring in either $S$ or $-S^{R}$ when $S$ is a $\mathcal{NOS}_{q}(n)$ and $\mathbf{v}$ is a negasymmetric $(n-1)$-tuple.
@@
[…]
It is important to observe that, from the definition of negative orientable, an element of $L^{*}_{S}(\mathbf{v})$, as with any $n$-tuple, can only appear at most once in $S$ or $-S^{R}$. We now state a result on which our bound is built.
@@
    Suppose $L^{*}_{S}(\mathbf{v})$ is non-empty. Then, by definition, and (without loss of generality) assuming that an element of $L(\mathbf{v})$ occurs in $S$ (as opposed to $-S^{R}$), we know that $v_{i}=s_{j+i}$ for some $j$ ($20\leq i\leq n-$ 2). Since $\mathbf{v}$ is negasymmetric, it follows immediately that $v_{i}=-s_{j+n-2-i}$ ($20\leq i\leq n-$ 2). That is, for each occurrence of an element of $L(\mathbf{v})$ in $S$, there is an occurrence of a, necessarily distinct, element of $L(\mathbf{v})$ in $-S^{R}$, and vice versa. The result follows.
If $|L(\mathbf{v})|$ is odd, Lemma 3.9 immediately shows that $S$ and $-S^{R}$ combined must omit at least one of the $n$-tuples in $L(\mathbf{v})$. We can now state our main result.
@@
  • First suppose that $q$ is odd. Suppose also that $\mathbf{v}$ is a negasymmetric non-uniform $(n-1)$-tuple. From Lemma 3.1 there are $q^{\lfloor(n-1)/2\rfloor}$ negasymmetric $(n-1)$-tuples of which precisely one, namely the all-zero tuple, is uniform — since $q$ is odd; i.e. there are $q^{\lfloor(n-1)/2\rfloor}-1$ such $\mathbf{v}$. From Lemma 3.5, none of the $q$ elements of $L(\mathbf{v})$ are negasymmetric, and hence $|L(\mathbf{v})\cap N^{*}_{q}(n)|=q$. Since $|L(\mathbf{v})|=q$ is odd, Lemma 3.9 implies that $L^{*}_{S}(\mathbf{v})$ must omit at least one of the $n$-tuples in $L(\mathbf{v})$. That is, there are $q^{\lfloor(n-1)/2\rfloor}-1$ non-negasymmetric $n$-tuples which cannot appear in $S$ or $-S$. The bound for $q$ odd follows from Corollary 3.2.
@@
We note that, as in the case of orientable sequences, all these bounds on negative orientable sequences are asymptotically of the same order as $q^{n}/2$.
@@ 3.2 Construction I: Sequences with $n=2$
The first method of construction we present provides negative orientable sequences of maximal length for every $q>3$ in the case $n=2$. The method involves joining certain explicitly defined cycles.
Consider the directed graph $G_{q}$ with vertex set $\mathbb{Z}_{q}$, $q\geq 3$, and all directed arcs $(x,y)$, $x,y\in\mathbb{Z}_{q}$, except for $(x,-x)$, $x\in\mathbb{Z}_{q}$. Then a $\mathcal{NOS}_{q}(2)$ $S$ corresponds to a circuit in $G_{q}$. Moreover $-S^{R}$ is also such a circuit and these two circuits are arc-disjoint. We say that a circuit $C=[c_{0},c_{1},\dots,c_{m-1}]$ of length $m$, consisting of arcs $(c_{i},c_{i+1})$, $i=0,\dots,m-1$, identifying $c_{m}$ with $c_{0}$, in $G_{q}$ is nice if it is arc-disjoint from $-C^{R}$. We note that if $C^{\prime}$ is another nice circuit in $G_{q}$ that is arc-disjoint from $C$ and $-C^{R}$ and if it shares a vertex with $C$ then $C$ and $C^{\prime}$ may be joined to make a nice circuit of length the sum of the lengths of $C$ and $C^{\prime}$.
@@
$C_{0}$ $=\left[0110220\dots 0\frac{q-1}{2}\frac{q-1}{2}\right]\text{~{}~{}of length~{}}3\frac{q-1}{2}$
$C_{1}$ $=\left[121(-2)131(-3)1\dots 1\frac{q-1}{2}1(-\frac{q-1}{2})\right]\text{~{}~{}of length~{}}4\frac{q-3}{2}$
@@
$C_{i}$ $=\left[i(i+1)i-(i+1)i\dots i\frac{q-1}{2}i-\frac{q-1}{2}\right]\text{~{}~{}of length~{}}4\frac{q-i-2}{2}$
@@
$C_{\frac{q-3}{2}}$ $=\left[\frac{q-3}{2}\frac{q-1}{2}\frac{q-3}{2}-\frac{q-1}{2}\right]\text{~{}~{}of length 4}.$
These circuits are arc-disjoint since by construction the arcs of $C_{i}$ involve the vertex $i$, with the exception of the loops $(j,j)$ in $C_{0}$, in such a way that there is no repetition. Moreover they are arc-disjoint to $-C_{i}^{R}$, $i=0,\dots,\frac{q-3}{2}$ since by construction the arcs of $-C_{i}^{R}$ involve the vertices $-j$ in such a way that there is no repetition. Thus they are mutually nice circuits and it may be seen that these circuits may be joined to give a $\mathcal{NOS}_{q}(2)$ $C$. A simple calculation shows that $C$ has period $\frac{q(q-1)}{2}$ and therefore is maximal by Theorem 3.10.
@@
$C_{0}$ $=\left[0110220\dots 0\frac{q-2}{2}\frac{q-2}{2}\right]\text{~{}~{}of length~{}}3\frac{q-2}{2}$
$C_{1}$ $=\left[121(-2)131(-3)1\dots 1\frac{q-2}{2}1(-\frac{q-2}{2})1\frac{q}{2}\right]\text{~{}~{}of length~{}}4\frac{q-4}{2}+2$
@@
$C_{i}$ $=\left[i(i+1)i-(i+1)i\dots i\frac{q-2}{2}i-\frac{q-2}{2}i\frac{q}{2}\right]\text{~{}~{}of length~{}}4\frac{q-2i-2}{2}+2$
@@
[…]